# Stable Nonconvex-Nonconcave Training via Linear Interpolation

### v1

@@ Abstract
This paper presents a theoretical analysis of linear interpolation as a principled method for stabilizing (large-scale) neural network training. We argue that instabilities in the optimization process are often caused by the nonmonotonicity of the loss landscape and show how linear interpolation can help by leveraging the theory of nonexpansive operators. We construct a new optimization scheme called relaxed approximate proximal point (RAPP), which is the first explicit method to achieve last iterate convergence rates for the full range of cohypomonotone problems. The construction extends to constrained and regularized settings. By replacing the inner optimizer in RAPP we rediscover the family of Lookahead algorithms for which we establish convergence in cohypomonotone problems even when the base optimizer is taken to be gradient descent ascent. The range of cohypomonotone problems in which Lookahead converges is further expanded by exploiting that Lookahead inherits the properties of the base optimizer. We corroborate the results with experiments on generative adversarial networks which demonstrates the benefits of the linear interpolation present in both RAPP and Lookahead.
@@ 1 Introduction
Table 1: Overview of last iterate results with our contribution highlighted in blue. Prior to this work there existed no explicit scheme handling $\rho\in(-\nicefrac{{1}}{{2L}},\infty)$ and no global convergence guarantees for Lookahead beyond bilinear games.
@@
A principled approach to cohypomonotone problems is the extragradient+ algorithm (EG+) proposed by Diakonikolas et al. (2021). However, the only known rates are on the best iterate, which can be problematic to pick in practice. It is unclear whether last iterate rates for EG+ are possible even in the monotone case (see discussion prior to Thm. 3.3 in Gorbunov et al. (2022a)). For this reason, the community has instead resorted to showing last iterate of extragradient (EG) method of Korpelevich (1977), despite originally being developed for the monotone case. Maybe not surprisingly, EG only enjoys a last iterate guarantee under mild form of cohypomonotonicity and have so far only been studied in the unconstrained case (Luo & Tran-Dinh, ; Gorbunov et al., 2022b). Recently it was shown that the full range of cohypomonotone problems is achievable, but the scheme is implicit and the complexity blows up with increasing cohypomonotonicity (Gorbunov et al., 2022b). This leaves the questions: Can an explicit scheme enjoy last iterate rates for all cohypomonotone problems? Can the rate be agnostic to the degree of cohypomonotonicity? We answer both in the affirmative.
As will become clear, the main mechanism behind convergence in this nonmonotone class is the linear interpolation also used in Lookahead. This iterative interpolation is more broadly referred to as the Krasnosel’skiĭ-Mann (KM) iteration in the theory of nonexpansive operators. We show that the extragradient+ algorithm (EG+) of Diakonikolas et al. (2021), our proposed relaxed approximate proximal point method (RAPP), and Lookahead based algorithms are all instances of the (inexact) KM iteration and provide simple proofs of these schemes in the cohypomonotone case.
@@
    We prove global convergence rates for the last iterate of our proposed algorithm RAPP which additionally handles constrained and regularized settings. This makes RAPP the first explicit scheme to have non-asymptotic guarantees for the full range of cohypomonotone problems. As a byproduct we obtain a last iterate convergence rate for an implicit scheme that is independent of the degree of cohypomonotonicity. The last iterate rates are established by showing monotonic decrease of the operator norm–something which is not possible for EG+. This contrast is maybe surprising, since RAPP can be viewed as an extension of EG+, which simply takes multiple extrapolation steps.
@@ 3 Setup
Figure 1: Consider $\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\phi(z)$ with $z=(x,y)$. As opposed to convex-concave minimax problems, the cohypomonotone condition allows the gradients $Fz=(\nabla_{x}\phi(z),-\nabla_{y}\phi(z))$ to point away from the solutions (see LABEL:app:weakMVI-comonotone for the relationship between cohypomonotonicity and the weak Minty variational inequality). This can lead to instability issues for standard algorithms such as the Adam optimizer.
@@
The problem can be recast as the inclusion problem (1) by defining the joint iterates $z=(x,y)$, the stacked gradients $Fz=(\nabla_{x}\phi(x,y),-\nabla_{y}\phi(x,y))$, and $A=(\mathcal{N}_{\mathcal{X}},\mathcal{N}_{\mathcal{Y}})$ where $\mathcal{N}$ denotes the normal cone. As will become clear (cf. Algorithm 1), we will only be accessing $A$ through the resolvent $J_{\gamma A}:=(\operatorname{id}+\gamma A)^{-1}$ which reduces to the proximal operator. More specifically $J_{\gamma A}(z)=(\operatorname{proj}_{\mathcal{X}}(x),\operatorname{proj}_{\mathcal{Y}}(y))$.
@@
    $\|Fz-Fz^{\prime}\|\leq L\|z-z^{\prime}\|\quad\forall z,z^{\prime}\in\mathbb{R}^{d}.$ (3)
@@ 5 Approximating the resolvent
Algorithm 1 Approximate proximal step
1:$w^{0}=z\in\mathbb{R}^{d}$, $\gamma\in(0,\nicefrac{{1}}{{L}})$
2:
3:$\xi_{t}\sim\mathcal{P}$
4:$w^{t+1}=(\operatorname{id}+\gamma A)^{-1}(z-\gamma\hat{F}_{\sigma}(w^{t},\xi_{t}))$
5:$w^{\tau}$
@@
Suppose 3.2(i), 3.2(ii), LABEL: and 3.4. Given $z\in\mathbb{R}^{d}$, the iterates generated by Algorithm 1 with $\gamma\in(0,\nicefrac{{1}}{{L}})$ converges to a neighborhood linearly, i.e.,
$\mathbb{E}\big{[}\|w^{\tau}-J_{\gamma S}(z)\|^{2}\big{]}\leq(\gamma L)^{2\tau}\|w^{0}-w^{\star}\|^{2}+\tfrac{\gamma^{2}}{(1-\gamma L)^{2}}\sigma^{2}.$ (6)
The resulting update in Algorithm 1 is identical to GDA but crucially always steps from $z$. We use this as a subroutine in RAPP to get convergence under a cohypomonotone operator while only suffering a logarithmic factor in the rate.
@@
for appropriately chosen $\mu\in(0,\infty)$. Algorithm 1 can thus be interpreted as solving a particular regularized subproblem. Later we will drop this regularization to arrive at the Lookahead algorithm.
@@ 6 Last iterate under cohypomonotonicity
All that remains to get convergence of the explicit scheme in RAPP, is to expand and simplify the errors $\varepsilon_{k}(z)$ and $\delta_{k}(z)$ using the approximation of the resolvent analyzed in Lemma 5.1.
@@ 10 Acknowledgements
This work was supported by Google. This work was supported by Hasler Foundation Program: Hasler Responsible AI (project number 21043). This work was supported by the Swiss National Science Foundation (SNSF) under grant number 200021_205011. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n° 725594 - time-data).

### v2

@@ Abstract
This paper presents a theoretical analysis of linear interpolation as a principled method for stabilizing (large-scale) neural network training. We argue that instabilities in the optimization process are often caused by the nonmonotonicity of the loss landscape and show how linear interpolation can help by leveraging the theory of nonexpansive operators. We construct a new optimization scheme called relaxed approximate proximal point (RAPP), which is the first 1-SCLI method to achieve last iterate convergence rates for $\rho$-comonotone problems while only requiring $\rho>-\tfrac{1}{2L}$. The construction extends to constrained and regularized settings. By replacing the inner optimizer in RAPP we rediscover the family of Lookahead algorithms for which we establish convergence in cohypomonotone problems even when the base optimizer is taken to be gradient descent ascent. The range of cohypomonotone problems in which Lookahead converges is further expanded by exploiting that Lookahead inherits the properties of the base optimizer. We corroborate the results with experiments on generative adversarial networks which demonstrates the benefits of the linear interpolation present in both RAPP and Lookahead.
@@ 1 Introduction
Table 1: Overview of last iterate results with our contribution highlighted in blue. Prior to this work there existed no rates for 1-SCLI schemes handling $\rho$-comonotone problems with $\rho\in(-\nicefrac{{1}}{{2L}},\infty)$ and no global convergence guarantees for Lookahead beyond bilinear games.
@@
A principled approach to cohypomonotone problems is the extragradient+ algorithm (EG+) proposed by Diakonikolas et al. (2021). However, the only known rates are on the best iterate, which can be problematic to pick in practice. It is unclear whether last iterate rates for EG+ are possible even in the monotone case (see discussion prior to Thm. 3.3 in Gorbunov et al. (2022a)). For this reason, the community has instead resorted to showing last iterate of extragradient (EG) method of Korpelevich (1977), despite originally being developed for the monotone case. Maybe not surprisingly, EG only enjoys a last iterate guarantee under mild form of cohypomonotonicity and have so far only been studied in the unconstrained case (Luo & Tran-Dinh, ; Gorbunov et al., 2022b). Recently, last iterate rate were established for the same (tight) range of cohypomonotone problems for which EG+ has best iterate guarantees. However, the analyzed scheme is implicit and the complexity blows up with increasing cohypomonotonicity (Gorbunov et al., 2022b). This leaves the questions: Can an explicit scheme enjoy last iterate rates for the same range of cohypomonotone problems? Can the rate be agnostic to the degree of cohypomonotonicity? We answer both in the affirmative.
This work focuses on 1-SCLI schemes (Arjevani et al., 2015; Golowich et al., 2020), whose update rule only depends on the previous iterate in a time-invariant fashion. Another approach to establishing last iterate is Halpern-type methods with an explicit scheme developed in Lee & Kim (2021) for cohypomonotone problems and later extended to the constrained case in Cai et al. (2022) (c.f. LABEL:app:relatework).
As will become clear, a principled mechanism behind convergence in this nonmonotone class is the linear interpolation also used in Lookahead. This iterative interpolation is more broadly referred to as the Krasnosel’skiĭ-Mann (KM) iteration in the theory of nonexpansive operators. We show that the extragradient+ algorithm (EG+) of Diakonikolas et al. (2021), our proposed relaxed approximate proximal point method (RAPP), and Lookahead based algorithms are all instances of the (inexact) KM iteration and provide simple proofs of these schemes in the cohypomonotone case.
@@
    We prove global convergence rates for the last iterate of our proposed algorithm RAPP which additionally handles constrained and regularized settings. This makes RAPP the first 1-SCLI scheme to have non-asymptotic guarantees for $\rho$-comonotone problems while only requiring $\rho>-\nicefrac{{1}}{{2L}}$. As a byproduct we obtain a last iterate convergence rate for an implicit scheme that is independent of the degree of cohypomonotonicity. The last iterate rates are established by showing monotonic decrease of the operator norm–something which is not possible for EG+. This contrast is maybe surprising, since RAPP can be viewed as an extension of EG+, which simply takes multiple extrapolation steps.
@@ 3 Setup
Figure 1: Consider $\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\phi(z)$ with $z=(x,y)$. As opposed to convex-concave minimax problems, the cohypomonotone condition allows the gradients $Fz=(\nabla_{x}\phi(z),-\nabla_{y}\phi(z))$ to point away from the solutions (see LABEL:app:weakMVI-comonotone for the relationship between cohypomonotonicity and the weak MVI). This can lead to instability issues for standard algorithms such as the Adam optimizer.
@@
The problem can be recast as the inclusion problem (1) by defining the joint iterates $z=(x,y)$, the stacked gradients $Fz=(\nabla_{x}\phi(x,y),-\nabla_{y}\phi(x,y))$, and $A=(\mathcal{N}_{\mathcal{X}},\mathcal{N}_{\mathcal{Y}})$ where $\mathcal{N}$ denotes the normal cone. As will become clear (cf. Algorithm 1), $A$ will only be accessed through the resolvent $J_{\gamma A}:=(\operatorname{id}+\gamma A)^{-1}$ which reduces to the proximal operator. More specifically $J_{\gamma A}(z)=(\operatorname{proj}_{\mathcal{X}}(x),\operatorname{proj}_{\mathcal{Y}}(y))$.
@@
    $\|Fz-Fz^{\prime}\|\leq L\|z-z^{\prime}\|\quad\forall z,z^{\prime}\in\mathbb{R}^{d}.$
@@ 5 Approximating the resolvent
$w^{t+1}=(\operatorname{id}+\gamma A)^{-1}(z-\gamma\hat{F}_{\sigma}(w^{t},\xi_{t}))\quad\xi_{t}\sim\mathcal{P}$ (5)
@@
The resulting update in (5) is identical to GDA but crucially always steps from $z$. We use this as a subroutine in RAPP to get convergence under a cohypomonotone operator while only suffering a logarithmic factor in the rate.
@@
for appropriately chosen $\mu\in(0,\infty)$. (5) can thus be interpreted as solving a particular regularized subproblem. Later we will drop this regularization to arrive at the Lookahead algorithm.
@@ 6 Last iterate under cohypomonotonicity
All that remains to get convergence of the explicit scheme in RAPP, is to expand and simplify the errors $\varepsilon_{k}(z)$ and $\delta_{k}(z)$ using the approximation of the resolvent analyzed in LABEL:thm:banach.
@@ 10 Acknowledgements
Thanks to Leello Dadi and Stratis Skoulakis for helpful discussion. This work was supported by Google. This work was supported by Hasler Foundation Program: Hasler Responsible AI (project number 21043). This work was supported by the Swiss National Science Foundation (SNSF) under grant number 200021_205011. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n° 725594 - time-data).